OpenStudy (anonymous):

Derivative of f(x)=sin^2 ((pi x)/4). Professor skimmed this and did not give good clarification in his example, nor are there good examples in the text.

OpenStudy (anonymous):

Chain rule. derivative of sin^2(x)=2sin(x)cos(x)*dx so derivative of sin^2(pi*x/4)=2sin(pi*x/4)cos(pi*x/4)*(pi/4)

OpenStudy (anonymous):

Did we lose a 2 in the simplified solution?

OpenStudy (anonymous):

simplified \[\frac{ \pi }{ 2 }\sin(\frac{ \pi x }{ 4 })\cos(\frac{ \pi x}{4})\]

OpenStudy (anonymous):

That looks more like what I was getting :)

OpenStudy (anonymous):

Thanks so much for your help :)